OpenStudy (anonymous):

HELP FOR GOD'S SAKE PLEASE!! 'A standard 27 inch analog tv screen as an aspect ratio of 4:3. Use proportions and the Pythagorean theorem to calculate the lengths and widths of the screen in inches.

OpenStudy (anonymous):

pythagorean thearam is horrible :P and I love the way you started this question lmao!

OpenStudy (anonymous):

i know the answer my teacher just wants the work.

OpenStudy (anonymous):

Just set up your right triangles. For the analog TV, the hypotenuse is 27 in, the length (l) = 4x, and the width (w) = 3x. This sets up your 4:3 aspect ratio. For the digital TV, the hypotenuse is again 27 in, the length (l) = 16x, and the width = 9x. This will provide you with the x value for each TV so you can figure out the length and width of each.

OpenStudy (anonymous):

If the ratio is 4:3 for the sides, then when you consider the triangle formed by adding the diagonal, you have two sides 3 and 4, meaning the hypotenuse (diagonal) is what?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

An interesting triangle in which the hypotenuse is longer than the sum of the other sides. Do you know what the Pythagoran Theorem says?

OpenStudy (anonymous):

well the answer to the problem is: 21.6in. by 16.2in. but I have to show the work and I tried almost everything!

OpenStudy (anonymous):

Follow my method (: I'll let the rest of you all handle this. Peace! XD

OpenStudy (anonymous):

noooo please help!!!

OpenStudy (anonymous):

If you prefer, the sides are 3x and 4x, making the hypotenuse what?

OpenStudy (anonymous):

I guess you can make the hypotunese 27, atleast thats what the teacher said...

OpenStudy (anonymous):

These guys are way smarter than I am lol and besides at the end of tonight you'll have this solved ♥

OpenStudy (anonymous):

Let's start from the beginning. Do you know what the Pythagorean Theorem says?

OpenStudy (anonymous):

that kid @iCanHelp gives smart answers though lol

OpenStudy (anonymous):

yes the pythagorean theorem says: a2 + b2 = c2

OpenStudy (anonymous):

On the TV screen, what are a, b, and c?

OpenStudy (anonymous):

a^2 + b^2 = c^2 (3x)^2 + (4x)^2 = 27^2 = 729 (3x)(3x) + (4x)(4x) = 729 9x^2 + 16x^2 = 729 25x^2 = 729 Does this help point you in the right direction?

OpenStudy (anonymous):

thats the problem! I only know that you have to use proportions AND the pythagoream theorem to figure this out

OpenStudy (anonymous):

but every method you guys are giving me I've already used, and it doesn't me the answer that is 21.6in by 16.2in

OpenStudy (anonymous):

If you draw a triangle consisting of a height, width, and the diagonal, you have a right triangle, OK? The screen is a rectangle. Of these three, which is the hypotenuse?

OpenStudy (anonymous):

Can i just post the whole EXACT problem to you guys so you can understand what I mean?

OpenStudy (anonymous):

I understand what the problem means. The question is, can you see the right triangle and identify the hypotenuse out of it?

OpenStudy (anonymous):

I don't normally like doing this and I'm sorry robspooner you're doing an excellent job explaining but... a^2 + b^2 = c^2 (3x)^2 + (4x)^2 = 27^2 = 729 (3x)(3x) + (4x)(4x) = 729 9x^2 + 16x^2 = 729 25x^2 = 729 x = sqrt(729/25) = 5.4 3x = 3*5.4 = 16.2 4x = 4*5.4 = 21.6 Understand?

OpenStudy (anonymous):

Of course he doesn't understand. And if he still doesn't understand when he writes down the answer, what's the point here?

OpenStudy (anonymous):

oh mother flabber you rock so much omg thankyou so much it means so much to me you don't even know!! thankyou so much @few

OpenStudy (anonymous):

Don't just write my answer down, crazyary8. Take the time to listen to what robspooner is saying and understand. If you're getting asked these questions, they WILL appear on tests when OpenStudy isn't available. This is a basic concept of mathematics that MUST be learned and committed to memory.

OpenStudy (anonymous):

No, I'm really good at math..it's my strongest subject. This is the very FIRST time i have ever asked somebody desperately for math help. People always come to me. I can't just write down the answer, cause I have to explain it to my teacher as well and when I see how you did it I will figure it out. I'm not like that don't worry (:

OpenStudy (anonymous):

how did you get the 27x though on the second line

OpenStudy (anonymous):

27x...? Where?